Grade 11Mechanics

A scooterist is moving on a straight road with speed 10m/s.A bus ahead of scooterist by 32m,starts from rest with an acceleration 1m/s2. If both are moving in the same direction then the time after which the scooter overtakes the bus is ?

ayush
7 Years agoGrade 11

1 Answer

Bhavya
7 Years ago
Time of overtaking = distance/v1-v2
V1=10
V2=? Using the eqn V^2-0=2×1×32    we get v2=8   
Therefore  from first eqn 
32/10-8=32/2=16s